What is the magnitude of the force that a 25μC-charge exerts on a 10μC charge {\rm 8.5\, cm}8.5cm away? (Take k=9×109 N.m2/C2)


1
Expert's answer
2022-03-15T10:33:55-0400

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{25*10^{-9}*10*10^{-9}}{(0.085)^2}\\
=311*10^{-6}\:\rm N"